OpenStudy (anonymous):

Express x^3 - 2 / 2x^3 - x^2 in partial fractions.

OpenStudy (luigi0210):

Factor the denominator

OpenStudy (anonymous):

Can you present it in a work solution?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

This question is very tricky i cant get the 1/2 for the first term of the answer.

OpenStudy (luigi0210):

Alright, I'll try hold on.

OpenStudy (luigi0210):

Do you have the answer?

OpenStudy (anonymous):

Answer is x^3 - 2 / x^3 + 2x = 1/2 + 4 / x + 2/x square - 15/2(2x-1)

OpenStudy (luigi0210):

@dumbcow take this >.<

OpenStudy (dumbcow):

\[\frac{x^{3}-2}{2x^{3} -x^{2}} = \frac{x^{3} -2}{x^{2}(2x-1)} = \frac{Ax+B}{x^{2}}+\frac{C}{2x-1}\] find A,B,C

OpenStudy (anonymous):

I tried this method already but i just couldnt get the 1/2 as the final answer as mentioned.

OpenStudy (luigi0210):

Yea, I got some weird answer with that method too. Can't you make x, and x^2?

OpenStudy (dumbcow):

oh right sorry you need a "x^3" term \[\rightarrow \frac{Ax^{2} +Bx+C}{x^{2}} +\frac{D}{2x-1}\] \[(Ax^{2} +Bx+C)(2x-1) +Dx^{2} = x^{3} -2\]

OpenStudy (luigi0210):

You sure about that?

OpenStudy (dumbcow):

A = 1/2 -A +2B+D = 0 --> D = 1/2 -2B -B +C = 0 --> C =B C = 2 B=2 D = -7/2

OpenStudy (anonymous):

x^\[x^3 - 2 / 2x^3 - x^2 = 1/2 + 0.5 x^2 - 2 / x^2 (2x - 1) \]

OpenStudy (anonymous):

can you interpret this step?

OpenStudy (anonymous):

= 1/2 + A/x + B/x^2 + C/ (2x-1)

OpenStudy (anonymous):

I think C the answer given was 15

OpenStudy (anonymous):

-15 sorry*

OpenStudy (dumbcow):

yes it was -15/2 i made 1 mistake

OpenStudy (anonymous):

can you type out the whole solution if possible. Thank you

OpenStudy (dumbcow):

basically you multiply it out and then set the coefficients of each term equal to each other

OpenStudy (anonymous):

If i were to separate the denominotor into x and x^2 would it be possible to find 1/2 as my first term answer? Thanks.

OpenStudy (dumbcow):

yes that is "A" .... A = 1/2

OpenStudy (anonymous):

how to get the 1/2 ?

OpenStudy (dumbcow):

maybe if i write it this way \[\frac{Ax^{2} +Bx +C}{x^{2}} = A + \frac{B}{x}+\frac{C}{x^{2}}\] so once you know A=1/2 , B=4, C=2, D = -15/2 from above it becomes \[\frac{1}{2}+\frac{4}{x}+\frac{2}{x^{2}}-\frac{15}{2(2x-1)}\]

OpenStudy (anonymous):

great job! now i understand already :D

OpenStudy (dumbcow):

:)